If Mr. Wee had 13 apples and 15 oranges. He pulls 2 without looking, what is the probability that they will both be apples?
Found 2 solutions by checkley71, stanbon:
Answer by checkley71(8403)

13/28*12/27=156/756=.206 or 20.6% probability that 2 apples are picked.

Answer by stanbon(48568)

If Mr. Wee had 13 apples and 15 oranges. He pulls 2 without looking, what is the probability that they will both be apples?
-----------------
P(2 apples) = 13C2/28C2 = 70/378 = 0.185185
----------------
Cheers,
Stan H.